OpenStudy (anonymous):

what is the curvature of r=5sin(2θ)?

OpenStudy (amistre64):

this sounds sooo familiar ....

OpenStudy (amistre64):

|'x''|/|cube'|

OpenStudy (anonymous):

calc 3, has to do with finding double derivative or something

OpenStudy (anonymous):

armistre64, thats what I was trying, but idk what r' and r" are

OpenStudy (anonymous):

*first and second derivative

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (amistre64):

really? derivative of a sin function is prettu standard

OpenStudy (anonymous):

Hi everybody, can I manipulate right on r or I have to transfer to x, and y?

OpenStudy (amistre64):

\[k=\frac{|r'xr''|||}{|r'|^3}\]

OpenStudy (anonymous):

yes, but it's not r' and r" its r(t)' and r(t)"

OpenStudy (anonymous):

is r even function of t?

OpenStudy (amistre64):

lol, my mistake, i assumed that a single variable was implied

OpenStudy (anonymous):

not its not, I'm not sure how to make r into a vector r(t)

OpenStudy (amistre64):

r = 5 sin(2t) r' = 10 cos(2t) r'' = -20 sin(2t)

OpenStudy (amistre64):

hmmm, i see a function format as well

OpenStudy (anonymous):

I was trying to use x=rcostheta and y=rsintheta to get <5sin(2θ)cos(θ), 5sin(2θ)sin(θ), 0 > as my vector

OpenStudy (amistre64):

r = 5 sin(2t) sin (2t) = 2 sin(t) cos(t) r(t) = 10 x(t) y(t)

OpenStudy (amistre64):

\[r = \sqrt{x^2+y^2}\] \[\sqrt{x^2+y^2} = 10xy\]

OpenStudy (anonymous):

I see where you're going with this

OpenStudy (anonymous):

so x2+y2 = 100x2y2

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so then that equals r, then what do I use as my vector?

OpenStudy (anonymous):

oh, I differentiate r = 100x2y2

OpenStudy (amistre64):

r = 10xy, r^2 = 100x^2y^2 :) lets stick with r=10xy and do partials is what im thinking, right? its been awhile

OpenStudy (anonymous):

I will have to give that a try, thank you!

OpenStudy (amistre64):

ill review on a few sites to clear some cobwebs

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

when r is defined as a function, it looks like\[k=\frac{|f''(x)|}{(1+(f'(x))^2)^{3/2}}\]

OpenStudy (amistre64):

\[k=\frac{|-20sin(2t)|}{(1+(10cos(2t))^2)^{3/2}}\]

OpenStudy (amistre64):

i just remembered how to parametrize this stuff :) x = r cos(t) = 5sin(2t)cos(t) = 5sin(t) cos^2(t) y = r sin(t) = 5sin(2t)sin(t) = 5 sin^2(t) cos(t)

OpenStudy (amistre64):

therefore; r = 5sin(t) cos^2(t) [i] + 5 sin^2(t) cos(t) [j] r' = a little bit of work .... lol

OpenStudy (anonymous):

for X hos did you get a cos&2(t)?

OpenStudy (anonymous):

and why does the y component have a cos in it?

OpenStudy (anonymous):

oh! nvm I see, you used the double angle ID for sin, thanks for the help!

OpenStudy (amistre64):

youre welcome ... and if its easier to leave it as 5 sin(2t) then by all means work it ;)